OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -9 - 3t. Find the instantaneous velocity at t = 8 by finding the derivative.

OpenStudy (anonymous):

Derive -9-3t. Do you know how to do that?

OpenStudy (anonymous):

@galacticwavesXX not exactly ... instruct me ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

So you have s(t) = -9-3t the derivative would be s'(t) = -9-3t Which is d/dt because it is in respect of t. d/dt = nx^n-1

OpenStudy (anonymous):

You understand what I wrote so far, correct?

OpenStudy (anonymous):

yes carry on :) so far lol

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

Yea sorry open study on iPad is acting confused

OpenStudy (anonymous):

okay so how is it done ?

OpenStudy (anonymous):

to get the answer

OpenStudy (anonymous):

We know that -9 is a constant and the derivative of a constant is 0. Also the power at t is 1 so now we can apply the rule d/dt = nx^n-1 where n is 1 because it is the power at t Which would give you s'(t) = -3 at t = 8

OpenStudy (anonymous):

thats it ? wow THANKS SO MUCH!

OpenStudy (anonymous):

one more ?

OpenStudy (anonymous):

You welcome Ok

OpenStudy (anonymous):

Here is the second one :)

OpenStudy (anonymous):

Give me a min been a while since I done a problem like that

OpenStudy (anonymous):

The vertical asymptote x=1

OpenStudy (anonymous):

As x approaches 1 the limit goes to infinity

OpenStudy (anonymous):

Those are the answers for that question